Graph these coordinates and answer the questions: (you will also have to draw a $4 tax on top of S and, separately, a $6 price floor). Initial equilibrium is (4, $4)

$ S qty D qty
X Y Y
0 0 8
$2 2 6
$4 4 4
$6 6 2
$8 8 0

Question 1

With a price floor of $6, what is the total amount that consumers pay?

  • a. Under $5
  • b. Between $5 and $7
  • c. Between $7 and $10
  • d. Over $10

Question 2

Which scenario would the Producers prefer?

No tax or floor, $2 tax, or $6 price floor?

  • a. No tax or floor.
  • b. $4 tax
  • c. $6 price floor
  • d. We cannot tell from the data
Answer #1

Q1. Option a.

From the above graph we can observe that the consumers would pay $2 after the tax of $4 which is less than $5

Q2. Option a

The revenue with no tax, 4*4 = 16

The revenue with $4 tax, 6*2 = 12

The revenue with $6 tax, 6*1 = 6

The revenue with price floor of $6= 6*2 = 12

Hence producer would favor no tax